# Flow Matching at Scale: A Machine Learning Framework for Efficient Large-Size Sampling of Many-Body Systems

### v3

@@ D.2 Optimization of Neural Network Architecture
The entire model is optimized using ”torch.compile” to enhance computational efficiency [54]. For certain blocks (”ResnetBlock” and ”ResnetBlock (Attention)” in Fig. 17) in model is optimized using ”epilogue fusion” and ”max autotune” supported by PyTorch [54] to perform operator fusion automatically. Please refer to the code in the following link for the detailed structure of the neural network. (https://github.com/ToelUl/Flow-U-Net)
@@ Appendix E Comparison between U-Net and DiT Architectures
To show that the inductive biases embedded in CNN architectures of U-Net are the key point for training on data of a fixed system size and extrapolating to other sizes, we compare the results using another robust neural network architecture: the Diffusion Transformer [37].
Vision Transformers [55] and, by extension, the Diffusion Transformer (DiT) [37] variants recently adopted for generative modeling, rely on minimal inductive biases. Transformer’s core is the self-attention mechanism [52]. This mechanism empowers them with global receptive fields and permutation equivariance, enabling each image patch (or token) to interact with all others regardless of spatial proximity. The only explicit inductive bias is introduced through the positional encoding scheme, which merely injects order information rather than enforcing locality. As a consequence, Transformers excel when long-range correlations or nonlocal dependencies dominate the target distribution, conditions frequently encountered in natural language, graphs, and physical systems near criticality where correlation lengths diverge. But in our case, the underlying data-generating process is governed by a physical Hamiltonian with strictly local interactions.
@@
This observation is consistent with a large body of literature in computer vision comparing Transformers and CNNs. The seminal work on Vision Transformers by Dosovitskiy et al. [55] explicitly states that Transformers lack the inductive biases of locality and translation equivariance, making them heavily reliant on massive-scale pre-training to generalize effectively. Subsequent analyses, such as the work by Raghu et al. [56], have provided further evidence, showing that ViTs must learn to represent local information from scratch, a task for which CNNs are architecturally predisposed. Our comparative experiment with the DiT model, therefore, serves as a direct demonstration of this principle within a physical simulation context: the model with the correct inductive bias (U-Net) not only performs better but does so with a fraction of the parameters, because its architecture is already aligned with the fundamental, local nature of the physical system.
@@ Appendix F GPU-accelerated Monte Carlo Method
We employ a highly optimized MCMC simulation framework implemented in PyTorch [54]. While the Wolff [7] or Swendsen–Wang [57] cluster algorithm is known to be more efficient in reducing autocorrelation steps for the 2D XY model, its sequential cluster-building process is not amenable to efficient GPU parallelization. Our choice of the Metropolis-Hastings algorithm with a checkerboard update scheme is motivated by its exceptional performance on massively parallel architectures, making it a powerful and relevant baseline for modern hardware. The full implementation of our algorithm is available on GitHub: (https://github.com/ToelUl/Thermal-Phase-Transition-MC-GPU-Simulation). The algorithm is designed to leverage the massively parallel architecture of modern Graphics Processing Units (GPUs), enabling large-scale simulations with high statistical precision. The core design incorporates three key parallelization strategies: a checkerboard update scheme for intra-lattice parallelism, the simulation of multiple independent Markov chains, and the replica-exchange method (also known as Parallel Tempering [10]) across a range of temperatures.
@@ F.2 Checkerboard Metropolis-Hastings Update
A standard single-spin flip Metropolis-Hastings algorithm introduces data dependencies, as the update of a spin requires knowledge of its neighbors, precluding simultaneous updates of adjacent spins. To overcome this limitation and maximize parallel efficiency, we implement a checkerboard update scheme [58]. The lattice is decomposed into two independent sub-lattices, ’A’ and ’B’ (analogous to the black and white squares of a checkerboard).
@@ F.3 Parallel Tempering (Replica-Exchange)
To mitigate critical slowing down near phase transitions and to ensure efficient exploration of the configuration space, our framework implements the Parallel Tempering (or replica-exchange) algorithm [10]. This method involves running simulations at multiple temperatures $T_{1},T_{2},\dots,T_{B}$ in parallel. Periodically, an exchange of the entire spin configurations between two adjacent replicas, say at temperatures $T_{i}$ and $T_{i+1}$ with energies $E_{i}$ and $E_{i+1}$, is proposed.
@@
  • Core Libraries [54, 59, 28]:

### v4

@@ D.2 Optimization of Neural Network Architecture
The entire model is optimized using ”torch.compile” to enhance computational efficiency [pytorch]. For certain blocks (”ResnetBlock” and ”ResnetBlock (Attention)” in Fig. 17) in model is optimized using ”epilogue fusion” and ”max autotune” supported by PyTorch [pytorch] to perform operator fusion automatically. Please refer to the code in the following link for the detailed structure of the neural network. (https://github.com/ToelUl/Flow-U-Net)
@@ Appendix E Comparison between U-Net and DiT Architectures
To show that the inductive biases embedded in CNN architectures of U-Net are the key point for training on data of a fixed system size and extrapolating to other sizes, we compare the results using another robust neural network architecture: the Diffusion Transformer [peebles2023DiT].
Vision Transformers [vit] and, by extension, the Diffusion Transformer (DiT) [peebles2023DiT] variants recently adopted for generative modeling, rely on minimal inductive biases. Transformer’s core is the self-attention mechanism [vaswani2023attention_need]. This mechanism empowers them with global receptive fields and permutation equivariance, enabling each image patch (or token) to interact with all others regardless of spatial proximity. The only explicit inductive bias is introduced through the positional encoding scheme, which provides spatial context but does not enforce strict locality. As a consequence, Transformers excel when long-range correlations or nonlocal dependencies dominate the target distribution, conditions frequently encountered in natural language, graphs, and physical systems near criticality where correlation lengths diverge. But in our case, the underlying data-generating process is governed by a physical Hamiltonian with strictly local interactions.
@@
This observation is consistent with a large body of literature in computer vision comparing Transformers and CNNs. The seminal work on Vision Transformers by Dosovitskiy et al. [vit] explicitly states that Transformers lack the inductive biases of locality and translation equivariance, making them heavily reliant on massive-scale pre-training to generalize effectively. Subsequent analyses, such as the work by Raghu et al. [vit_vs_cnn], have provided further evidence, showing that ViTs must learn to represent local information from scratch, a task for which CNNs are architecturally predisposed. Our comparative experiment with the DiT model, therefore, serves as a direct demonstration of this principle within a physical simulation context: the model with the correct inductive bias (U-Net) not only performs better but does so with a fraction of the parameters, because its architecture is already aligned with the fundamental, local nature of the physical system.
@@ Appendix F GPU-accelerated Monte Carlo Method
We employ a highly optimized MCMC simulation framework implemented in PyTorch [pytorch]. While the Wolff [PhysRevLett_Wolff] or Swendsen–Wang [PhysRevLett.58.86] cluster algorithm is known to be more efficient in reducing autocorrelation steps for the 2D XY model, its sequential cluster-building process is not amenable to efficient GPU parallelization. Our choice of the Metropolis-Hastings algorithm with a checkerboard update scheme is motivated by its exceptional performance on massively parallel architectures, making it a powerful and relevant baseline for modern hardware. The full implementation of our algorithm is available on GitHub: (https://github.com/ToelUl/Thermal-Phase-Transition-MC-GPU-Simulation). The algorithm is designed to leverage the massively parallel architecture of modern Graphics Processing Units (GPUs), enabling large-scale simulations with high statistical precision. The core design incorporates three key parallelization strategies: a checkerboard update scheme for intra-lattice parallelism, the simulation of multiple independent Markov chains, and the replica-exchange method (also known as Parallel Tempering [Hukushima_1996]) across a range of temperatures.
@@ F.2 Checkerboard Metropolis-Hastings Update
A standard single-spin flip Metropolis-Hastings algorithm introduces data dependencies, as the update of a spin requires knowledge of its neighbors, precluding simultaneous updates of adjacent spins. To overcome this limitation and maximize parallel efficiency, we implement a checkerboard update scheme [Weigel_2011]. The lattice is decomposed into two independent sub-lattices, ’A’ and ’B’ (analogous to the black and white squares of a checkerboard).
@@ F.3 Parallel Tempering (Replica-Exchange)
To mitigate critical slowing down near phase transitions and to ensure efficient exploration of the configuration space, our framework implements the Parallel Tempering (or replica-exchange) algorithm [Hukushima_1996]. This method involves running simulations at multiple temperatures $T_{1},T_{2},\dots,T_{B}$ in parallel. Periodically, an exchange of the entire spin configurations between two adjacent replicas, say at temperatures $T_{i}$ and $T_{i+1}$ with energies $E_{i}$ and $E_{i+1}$, is proposed.
@@
  • Core Libraries [pytorch, cudnn, lipman2024flow_matching_guide_code]: